### v1

# Knowledge-based Consistency Testing of Large Language Models

Abstract
In this work, we systematically expose and measure the inconsistency and knowledge gaps of Large Language Models (LLMs). Specifically, we propose an automated testing framework (called KonTest) which leverages a knowledge graph to construct test cases. KonTest probes and measures the inconsistencies in the LLM’s knowledge of the world via a combination of semantically-equivalent queries and test oracles (metamorphic or ontological oracle). KonTest further mitigates knowledge gaps via a weighted LLM model ensemble. Using four state-of-the-art LLMs (Falcon, Gemini, GPT3.5, and Llama2), we show that KonTest generates 19.2% error inducing inputs (1917 errors from 9983 test inputs). It also reveals a 16.5% knowledge gap across all tested LLMs. KonTest’s mitigation method reduces LLM knowledge gap by 32.48%. Our ablation study further shows that GPT3.5 is not suitable for knowledge-based consistency testing because it is only 60%-68% effective in knowledge construction.
Knowledge-based Consistency Testing of Large Language Models
Sai Sathiesh Rajan SUTD, Singapore sai_rajan@mymail.sutd.edu.sg                        Ezekiel Soremekun Royal Holloway, University of London ezekiel.soremekun@rhul.ac.uk
Sudipta Chattopadhyay SUTD, Singapore sudipta_chattopadhyay@sutd.edu.sg
1 Introduction
Large language models (LLMs) are being increasingly utilized in real-world applications. LLMs are powerful in solving many tasks, but their reliability remains a concern Qiu et al. (2023). This is alarming because inconsistent behaviors adversely affect critical downstream tasks and influence adoption.
In this paper, we study the problem of assessing inconsistency in LLM behaviors. Previous works have demonstrated the prevalence and severity of inconsistent responses in LLMs Min et al. (2023); Berglund et al. (2023); Sallou et al. (2023). To address this challenge, we conceptualize and design KonTest  [^1] KonTest means Knowledge-based Consistency Testing – a novel test generation methodology to systematically discover consistency errors in LLMs and highlight their knowledge gaps.
Figure 1 shows examples of consistency errors and knowledge gap discovered by KonTest in GPT3.5. These are errors in GPT3.5 about the place (spatial) domain. These errors may have adverse effects in critical areas (e.g., automobiles, aeronautics) where spatial LLMs are deployed, e.g., navigation systems – MapBox’s MapGPT Mapbox , LLM-Geo Li and Ning (2023) and MapGPT Chen et al. (2024), L3MVN Yu et al. (2023)). KonTest allows to automatically discover and expose such errors to end-users/developers. This is the first step to enable their mitigation and repair for LLM improvement.
Figure 1: Examples of queries generated and errors discovered by KonTest in GPT3.5: Error reflected (a) in two different conversations asking semantically equivalent questions (Atomic error in Table 1) and (b) in conversation asking semantically equivalent questions in a sequence (Sequential error in Table 1). (c) Example of (Ontological) error in Table 1, (d) the knowledge graph extracted for the considered entities i.e., Ulster, Kinawley and Ireland, (e) SUT’s knowledge showing that the LLM lacks the knowledge about Kinawley is in Ireland.
KonTest leverages a knowledge graph for consistency testing (see Figure 2). It first automatically extracts a set of entities and entity relationships from the knowledge graph. This is then used to systematically generate (semantically-equivalent) queries and create test cases for the subject LLM under test (SUT) across various settings (e.g., atomic query vs. sequential query). The test case generation involves minimal, one-time effort for each type of entity relation considered (e.g., only two templates for 6730 test cases) and such templates can be reused for testing arbitrary LLMs. Additionally, KonTest automatically mitigates knowledge gaps by leveraging the likelihood of LLM inconsistencies to construct a weighted model ensemble. To the best of our knowledge, KonTest is the first systematic approach for automatically generating consistency tests for assessing LLMs.
Knowledge-based test generation provides a unique and systematic method for exploring the SUT’s knowledge and compute a test adequacy metric for the SUT in terms of the covered entities and relations. More importantly, the responses from the SUT allow KonTest to construct the SUT’s knowledge base as a subset of the extracted knowledge graph and to concretely highlight the knowledge gap of the SUT. This is valuable information for developers, it enables refining the model and improving its knowledge.
KonTest differs from existing works both in its objective and in its unique approach for testing LLMs. Prior works have focused on other non-functional properties such as reasoning ability, robustness and security Honarvar et al. (2023); Wu et al. (2023); Yang et al. (2023b). Few works have demonstrated the importance of LLM consistency, e.g., via self-consistency Min et al. (2023) and reversal curse Berglund et al. (2023). In contrast, KonTest employs automated test generation to discover a larger scope of consistency errors.
This paper provides an overview of KonTest (section 2), and makes the following contributions:
  1. 1.
    We present KonTest, a novel approach to leverage knowledge graph for checking the consistency of LLM responses and measure their knowledge gaps (section 3).
  2. 2.
    We present the metamorphic and ontological oracles that allow to check consistency errors in LLM results (section 3).
  3. 3.
    We implement KonTest and evaluate it with Falcon Nomic AI (2023), Gemini Anil et al. (2023), GPT3.5 OpenAI (2023), and Llama2 Touvron et al. (2023). KonTest revealed 19.2% erroneous inputs and exposed an average knowledge gap of 16.5% (section 5).
  4. 4.
    KonTest automatically mitigates knowledge gaps via a weighted model ensemble via the likelihood of LLM inconsistencies. The mitigation technique of KonTest reduces knowledge gaps by 32.48% (section 5).
  5. 5.
    We perform an ablation study by replacing each component of KonTest with GPT3.5 using few-shot prompting. We found that GPT3.5 constructs at most 68% of the ground truth knowledge base. It exhibits up to 63.7% false positives in error detection (section 5).
After the related works (section 6), we conclude in section 7 and discuss limitations (section 8).
2 Overview
In this section, we outline the motivation behind our approach and present an illustrative example to demonstrate the overall process of our approach.
Figure 2: Overall workflow of our approach (KonTest)
Table 1: Test cases generated and error types detected by KonTest using metamorphic and ontological oracles.
Generated Sentences Error Type Oracle
atomic
LLM(Q1)=LLM(Q2)
sequential-intra
LLM(Q5a)=LLM(Q5b),
LLM(Q6a)=LLM(Q6b)
sequential-inter
LLM(Q1)=LLM(Q5b),
LLM(Q2)=LLM(Q6b)
Ontological
(LLM(Q3)=LLM(Q4)=Yes) $\rightarrow$
(LLM(Q2)=Yes)
Key Insight: The key insight behind KonTest is to leverage a knowledge graph for systematic consistency testing of LLMs. The knowledge graph serves multiple purposes in KonTest: Firstly, entities and entity relations extracted from the knowledge graph allows KonTest to systematically generate queries and construct test cases for validating the consistency of the subject LLM (SUT). The LLM’s responses to these test cases allow KonTest to build the knowledge (sub)-graph where the LLM behaves correctly. Secondly, knowledge-based test generation allows KonTest to report a test adequacy metric for the SUT in terms of the entities and relations covered within the knowledge graph. Finally, this approach enables KonTest to highlight sub-graph of the knowledge graph where the outputs from the SUT are logically inconsistent. This enables practitioners to selectively focus on the knowledge gaps highlighted by KonTest for improving the LLM (e.g., by fine-tuning).
Figure 3: Example queries and templates in KonTest
[…]
3 Methodology
3.1 Knowledge Graph based Test Generation
[…]
Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path. To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes. KonTest then generates a query, ${Query}$, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (9 in Algorithm 1). KonTest also generates the mutated query, ${Query_{M}}$, with another template (12 in Algorithm 1). These templates are dependent on the relation between the two nodes in question. For instance, Figure 3 shows the set of queries and mutated queries (i.e., $Query_{list}$ and $MutQuery_{list}$) generated from one particular path and along with the respective templates used to generate them. We note that developers can implement additional templates easily if they are interested in querying different relations. Moreover, creation of these templates is a one-time process and the cost incurred is minimal since the number of different types of relationships is small in comparison to the number of entities.
Algorithm 1 KG-Based Query Generation
1:procedure gen_test($\mathbb{KG}$)
2:     $Query_{list},MutQuery_{list}\leftarrow\varnothing$
3:     for $\mathbb{KG}_{node1}$ in $\mathbb{KG}$ do
4:         for $\mathbb{KG}_{node2}$ in $\mathbb{KG}$ do
5:              if $\mathbb{KG}_{node1}\neq\mathbb{KG}_{node2}$ then
6:                  $\triangleright$ Finds relationship between $\mathbb{KG}_{node1}$ and $\mathbb{KG}_{node2}$
7:                  $\mathbb{KG}_{rel}\leftarrow$ Find_Relation($\mathbb{KG}_{node1},\mathbb{KG}_{node2}$)
8:                  $\triangleright$ Builds query using the relation ($\mathbb{KG}_{node1}$, $\mathbb{KG}_{node2}$)
9:                  $Query\leftarrow$ Gen_Query($\mathbb{KG}_{node1},\mathbb{KG}_{node2},\mathbb{KG}_{rel}$)
10:                  $Query_{list}\leftarrow Query_{list}\cup\{{Query}\}$
11:                  $\triangleright$ Generates semantically equivalent query
12:                  ${Query_{M}}\leftarrow$ Mut_Gen(${Query}$)
13:                  $MutQuery_{list}\leftarrow MutQuery_{list}\cup\{{Query_{M}}\}$
14:     return $Query_{list},MutQuery_{list}$
3.2 Test Oracles
KonTest detects consistency errors in LLM via a metamorphic oracle and an ontological oracle.
Metamorphic Oracle: This oracle leverages both the original ($Query_{list}$) and mutated ($MutQuery_{list}$) query sets to create four unique conversations. These conversations are then fed into the target LLM (SUT). Concretely, KonTest creates two conversations containing the answer to both (initial and mutated) atomic queries individually (see the first column of “Generated Sentences" in Table 1) and two conversations where the queries are fed in a sequential manner (see the second column of “Generated Sentences" in Table 1). Responses from these conversations are then evaluated using the consistency checker embodied in KonTest to identify the erroneous test cases and the number of each error types (i.e., atomic, sequential-intra and sequential-inter in Table 1).
Ontological Oracle: Firstly, this oracle builds the SUT’s knowledge graph. To this end, the generated queries from KonTest are fed into the SUT and the target responses are recorded. Subsequently, given a list of such responses, KonTest identifies the nodes involved in each response. These nodes are then added to the SUT’s knowledge graph. Next, if the query response is positive (i.e., correct), KonTest introduces a directed edge between the nodes in the SUT’s knowledge graph indicating that the respective relation exists between the two nodes being considered. It is important to note that the relations in question are not symmetric in nature, justifying the directed flavor of the edge. After constructing the SUT’s knowledge graph, ${Graph_{Gen}}$, KonTest checks for ontological consistency errors with the aid of Algorithm 2.
Algorithm 2 Graph Checker
1:procedure Graph_Checker(${Graph_{Gen}}$)
2:     ${Err_{count}},{Coverage_{count}}\leftarrow\phi$
3:     for $Node_{1}$ in $Graph_{Gen}$ do
4:         for $Node_{2}$ in $Graph_{Gen}$ do
5:              if $Node_{1}\neq Node_{2}$ then
6:                  $\triangleright$ Checks whether direct path exists between nodes.
7:                  ${Path_{Dir}}\leftarrow$ Dir_Path(${Node_{1}},{Node_{2}}$)
8:                  if $Path_{Dir}\neq TRUE$ then
9:                       $\triangleright$ Checks whether indirect path exists between nodes.
10:                       ${Path_{InDir}}\leftarrow$ InDir_Path(${Node_{1}},{Node_{2}}$)
11:                       if $Path_{InDir}==TRUE$ then
12:                           ${Err_{count}}\leftarrow{Err_{count}}+1$
13:                  else
14:                       ${Coverage_{count}}\leftarrow{Coverage_{count}}+1$
15:     return $Err_{count},{Coverage_{count}}$
To check the consistency in SUT’s knowledge graph, Algorithm 2 first inspects whether a direct path exists between any pair of nodes ($Node_{1},Node_{2}$) in ${Graph_{Gen}}$ (7 in Algorithm 2). In the absence of such a direct path, when an indirect path exists between the same pair of nodes (11 in Algorithm 2), KonTest detects an ontological consistency error and updates the error count, ${Err_{count}}$. Intuitively, a direct path between two nodes indicates that the SUT responded positively to a question relating the two nodes. Similarly, the presence of an indirect path indicates that SUT knows that a positive relationship exists between the nodes albeit through more than one queries. As such, the SUT can be considered to be exhibiting an ontological inconsistency. As an example, consider the ontological oracle illustrated in Table 1. In this case, if an indirect path exists between nodes for Kinawley and Ireland via node Ulster (i.e., the responses to both Q3 and Q4 are positives from the SUT), then KonTest concludes that a direct path should also exist between Kinawley and Ireland. Hence, the absence of a positive response for the respective query (Q2) is indicated as an ontological error. As a byproduct of our ontological oracle, KonTest computes the knowledge coverage ${Coverage_{count}}$ for the SUT.
3.3 Knowledge Gap Mitigation
[…]
KonTest then uses the relative weights to calculate the final ensemble score by first multiplying the relative weights with the results for each response and summing them. In particular, it considers a positive and negative responses to be one and zero respectively. If the ensemble score is above the midpoint (0.5), the ensemble response for the relation is considered to be positive. To determine the effectiveness of our weighting, we compare it to simple majority voting – an ensemble that does not account for the likelihood of inconsistencies for each SUT (see RQ3).
4 Evaluation Setup
To evaluate our approach (KonTest), we pose the following research questions (RQs):
  • RQ1 Effectiveness: How effective is KonTest in exposing consistency errors in LLMs?
  • RQ2 Knowledge Coverage: What is the SUT’s knowledge graph coverage? How much is the SUT knowledge gap revealed by KonTest?
  • RQ3 Knowledge Gap Mitigation: How effective is KonTest’s weighted ensemble technique in improving knowledge coverage? How does it compare to a simple majority voting? Does it generalize to a new template?
  • RQ4 Ablation Study with GPT3.5: Can the state-of-the-art LLM (GPT3.5) perform the three main sub-tasks of KonTest, i.e., knowledge construction, test generation and error detection?
Knowledge Graph: We rely on Wikidata’s knowledge graph Vrandečić and Krötzsch (2014) and Wikidata’s SPARQL query service to extract information pertinent to the two tested domains: places and music. We choose these domains as inconsistent knowledge about places may have adversarial consequences in navigation use cases of LLMs Mapbox . Concurrently, inconsistencies in music domain may inaccurately capture the ownership of copyrighted materials (e.g., music albums). For the places domain, we take the list of countries ranked by (nominal) GDP per capita and select the top ten countries as our initial entity list IMF . We then iteratively build our knowledge graph by finding “administrative divisions" that are “located in the administrative territorial entity" of the parent entity. For the music domain, we take the top 200 musical acts since 2000 as our initial entity list Chart (2000). We then find the set of “albums" that have the parent entity as a “performer". We also find the set of songs and singles that are “part of" the album being considered. Once the graph is fully constructed, we randomly select 50 leaf nodes for each domain and extract the full path associated with the selected leaf nodes.
Subject Programs and Query Construction: Table 2 outlines features of the LLMs (SUTs) tested by KonTest. We ensure that temperature for each query is set to zero where possible. In addition, we provide the LLMs with an additional system command, “Be concise as possible. Answer with a yes or no response,", before the start of any conversation. In the event that the LLM does not support a system command, we prepend the statement to the queries before feeding it to the LLM. For the queries themselves, we construct two sets of queries with the aid of a template relating the parent and child entities, as illustrated in Figure 3.
Test Generation and Evaluation: We provide each LLM with each of the queries present in the two sets (i.e., original and mutated) in an atomic manner. We then feed the queries in a sequential manner with queries from both sets being presented in turn. This is then repeated with the order reversed. To invoke the oracles, we only consider responses that begin with a yes or no. Other responses are classified as “Invalid". This allows us to determine the accuracy of the responses without penalizing LLMs that fail to respond appropriately.
Table 2: Model Details
LLM Lineage Model Size Release Date
GPT3.5 GPT (OpenAI) gpt-3.5-turbo 175B Mar 2023
Falcon
Falcon (TII)
Finetuned by Nomic
gpt4all-falcon 7B Jun 2023
Llama2 Llama (Meta) Llama-2-7B 7B Jul 2023
Gemini Gemini (Google) gemini-pro Unknown Dec 2023
Knowledge Gap Mitigation: We evaluate the effectiveness of KonTest’s ensemble method by determining relative weights assigned to each SUT. We restrict the set of relations considered for this process such that a relation that yields an invalid answer in the query generation phase for any of the SUTs is not used in subsequent computations. We then perform five-fold cross validation on the remaining set of relations. Concretely, we hold out 20% of the relations for evaluation and use the consistency errors associated with the rest of the relations in the computation of the weights. We introduce a third template to evaluate whether KonTest’s mitigation generalizes to unseen templates.
Ablation Study: We assess whether an LLM could conceivably be used to replicate each of the sub-tasks in our approach by testing its applicability on the places domain. In particular, we chose to use GPT3.5 for this due to its popularity and relatively high percentage (99.9%) of non-exempted responses in our experiments. In each case described below, we provide GPT3.5 with sample questions and answers (few shot prompts) to ensure that the outputs adhere to the format of KonTest.
To test the knowledge construction capability of GPT3.5, we ask GPT3.5 where each of the 50 leaf nodes (used to evaluate KonTest) is located. We then check the accuracy of the response and we exempted responses that were not in the requested format. To evaluate GPT3.5’s test generation capability, we provide the relation we are concerned with and ask it to generate two semantically equivalent queries involving the two entities. For ontological oracle, we provide GPT3.5 with the full set of relations for each (knowledge) path and ask it to generate queries that expose inconsistencies in an LLM. For both oracles, if entity names and relations (as provided in the prompt) are not present in a generated query, then the test is considered invalid. For instance, “Is there a County of Capellen in County of Capellen?" is invalid since it does not correspond to the given knowledge relations. For error detection, we provide each set of queries along with their associated responses from each subject to GPT3.5 and ask it to classify each set as being Consistent or Inconsistent. We also allow it to respond with Invalid when the SUT responses were exempted. For the ontological oracle, we provide GPT3.5 with all the queries and responses relating to the entirety of a path and ask it to identify inconsistencies. We note that KonTest performs its check in Algorithm 2 using the same set of inputs provided to GPT3.5.
Implementation Details and Platforms: KonTest utilizes PyTorch 2.0, CUDA 11.3 and the llm package. All experiments were conducted on a GCP VM with an N1 series machine, eight vCPUs, 30 GB of memory and one Nvidia T4 GPU.
Table 3: Effectiveness of KonTest using a knowledge graph. (Meta.: Metamorphic, Onto.: Ontological)
LLM (Subject)
Valid Test
Executions (%)
Errors (%)
Meta. Onto. All Meta. Onto. All
Places Falcon 1322 (89.9) 267 (91.4) 1589 (90.2) 226 (17.1) 25 (9.4) 251 (15.8)
Gemini 1470 (100) 292 (100) 1762 (100) 161 (11.0) 40 (13.7) 201 (11.4)
GPT3.5 1468 (99.9) 292 (100) 1760 (99.9) 400 (27.2) 17 (5.8) 417 (23.7)
Llama2 1351 (91.9) 268 (91.8) 1619 (91.9) 355 (26.3) 27 (10.1) 382 (23.6)
Total 5611 (95.4) 1119 (95.8) 6730 (95.5) 1142 (20.4) 109 (9.7) 1251 (18.6)
Music Falcon 715 (97.9) 94 (97.9) 809 (97.9) 131 (18.3) 3 (3.2) 134 (16.6)
Gemini 730 (100) 96 (100) 826 (100) 124 (17.0) 7 (7.3) 131 (15.9)
GPT3.5 730 (100) 96 (100) 826 (100) 181 (24.8) 5 (5.2) 186 (22.5)
Llama2 696 (95.3) 96 (100) 792 (95.9) 206 (29.6) 9 (9.4) 215 (27.1)
Total 2871 (98.3) 382 (99.5) 3253 (98.5) 642 (22.4) 24 (6.3) 666 (20.5)
Total 8482 (96.4) 1501 (96.7) 9983 (96.4) 1784 (21.0) 133 (8.9) 1917 (19.2)
5 Evaluation Results
Table 4: Knowledge Gap Mitigation results for KonTest’s weighted ensemble (“KonTest”), versus simple majority voting ensemble (“Majority”) and the initial average knowledge gap found in the SUTs (“SUTs”). Best reduction in knowledge gaps are in bold text. “%Impr. ” means percentage improvement over the SUT.
Domain Relations Knowledge Gap (%)
Template 1 Template 2 Template 3 All (%Impr.)
KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs
Places 224 28 (12.5) 38 (17.0) 38.25 (17.1) 49 (21.9) 77 (34.4) 65.75 (29.4) 30 (13.4) 49 (21.9) 41.5 (18.5) 107 (26.46) 164 (-12.71) 145.5
Music 132 39 (29.5) 60 (45.5) 48.5 (36.7) 15 (11.4) 44 (33.3) 35.75 (27.1) 24 (18.2) 55 (41.7) 44.25 (33.5) 78 (39.30) 159 (-23.74) 128.5
Total 356 67 (18.8) 98 (27.5) 86.75 (24.4) 64 (18.0) 121 (34.0) 101.5 (28.5) 54 (15.2) 104 (29.2) 85.75 (24.1) 185 (32.48) 323 (-17.88) 274
RQ1 Effectiveness: We found that KonTest is effective in exposing consistency errors in LLMs: 19.2% of valid test executions result in consistency errors. Table 3 also shows that metamorphic errors (21.0%) are more prevalent than ontological errors (8.9%). We attribute the lower error rate of ontological errors/inputs to the high complexity of our ontological oracle (see Table 1). Table 3 demonstrates that metamorphic errors are common across all LLMs. In particular, the tested LLMs are highly inconsistent when asked the same question in a different manner (see Table 1). In addition, we observe that a large model size does not necessarily lead to a smaller error rate than a smaller model, e.g., GPT3.5 exhibits a metamorphic error rate of 27.2% for the places entities as compared to a metamorphic error rate of 17.1% for Falcon.
KonTest effectively exposes consistency errors in LLMS: 19.2% of valid test executions exposed a metamorphic or ontological error in LLMs.
RQ2 Knowledge Coverage: Table 5 shows that the tested LLMs cover about four-fifth (83.5%) of the tested knowledge graph across both templates. We found that LLMs are particularly sensitive to the query templates. For instance, Llama2 exhibits a 47.3% gap in knowledge for the places entities for one template, but only exhibits a 12.9% knowledge gap for the other. This suggests that LLMs may respond differently to two different, but semantically equivalent, input queries. We also observed that the smallest model (Falcon) has the lowest knowledge gap for the places entities (3.1%), while one of the biggest models (GPT3.5) has the highest knowledge gap (29.6%). This implies that model size/complexity is not a good proxy for knowledge coverage/gap. We do, however, attribute the performance of Falcon to its tendency to answer with a positive response regardless of the query. In addition, we also note that all queries posed to the subject LLMs are queries relating to an existing relation. These results show that KonTest effectively reveals the knowledge gap in LLMs. We posit that knowledge coverage is a good criteria for estimating the underlying knowledge of an LLM.
KonTest effectively reveals knowledge gaps in LLMs: It exposes an average knowledge gap of 16.5% in the tested LLMs.
Table 5: Knowledge Coverage and Gap in tested LLMs (Highest coverage or gap are marked in bold text)
LLM (Subject) Knowledge Gap (%) Overall Coverage (%)
Relations Template 1 Template 2 Intersection
Places Falcon 294 12 (4.1) 93 (31.6) 9 (3.1) 285 (96.9)
Gemini 294 60 (20.4) 71 (24.1) 40 (13.6) 254 (86.4)
GPT3.5 294 129 (43.9) 107 (36.4) 87 (29.6) 207 (70.4)
Llama2 294 38 (12.9) 139 (47.3) 37 (12.6) 257 (87.4)
Music Falcon 146 20 (13.7) 15 (10.3) 3 (2.1) 143 (97.9)
Gemini 146 60 (41.1) 45 (30.8) 33 (22.6) 113 (77.4)
GPT3.5 146 63 (43.2) 27 (18.5) 25 (17.1) 121 (82.9)
Llama2 146 76 (52.1) 84 (57.5) 56 (38.4) 90 (61.6)
Total 1760 458 (26.0) 581 (33.0) 290 (16.5) 1470 (83.5)
RQ3 Knowledge Gap Mitigation: Results show that KonTest’s mitigation technique reduces the SUT’s knowledge gap by up to 39.30%. Table 4 shows that KonTest reduces the knowledge gap for all templates by 32.48%. We found that the simple majority voting ensemble worsens the SUTs’ knowledge gap by up to 23.74%. We also observed that the mitigation performance of KonTest generalizes to an unseen template (template three (3)). While the mitigation performance of KonTest is slightly better on the new template than the original templates the simple majority performs worse on an unseen template. These results demonstrate the generalizability of KonTest’s mitigation and the efficacy of our weighted ensemble. We attribute the performance of KonTest’s mitigation to the likelihood and distribution of inconsistencies: Firstly, KonTest’s weighted ensemble is 42.72% more effective than the simple majority voting. Secondly, the distribution of inconsistencies shows that only 1.1% (15 out of 1330) of errors are found in all four SUTs while 65.7% (874 out of 1330) of errors are unique to a single SUT.
The ensemble method of KonTest effectively reduces the SUT’s knowledge gap by 32.48%.
RQ4 Ablation study with GPT3.5: We conduct an ablation study to investigate whether a state-of-the-art LLM (GPT3.5) is as effective as KonTest in performing its three main sub-tasks – knowledge construction, test generation and error detection.
Table 6: GPT3.5 efficacy in knowledge construction
Number of LLM Responses (%)
Total No Response Exempted Unexempted Correct Incorrect
Nodes 50 0 (0) 5 (10.0) 45 (90.0) 30 (60.0) 15 (30.0)
Edges 294 39 (13.3) 30 (10.2) 225 (76.5) 200 (68.0) 25 (8.5)
Knowledge Construction: Table 6 demonstrates that GPT3.5 is not a reliable knowledge constructor. GPT3.5 is only able to identify 68% of the relations present in the original knowledge graph. We believe this is because LLMs are generally intended to be a conversational engine, rather than a knowledge database Pan et al. (2023). These results show the importance of knowledge graphs in KonTest and suggests that LLMs are not a reliable replacement for knowledge graphs.
GPT3.5 is an ineffective surrogate for a knowledge graph: It constructs only 68% of the ground-truth entity relations (edges).
Test Generation: Given a few (three) query examples and the relations in a knowledge graph, GPT3.5 is slightly more effective (17.5% vs 16.8%) than KonTest in test generation (see Table 7). We also observe that this effectiveness persists for both the metamorphic (18.7% vs 17.9%) and the ontological oracle (11.5% vs 11.1%). We attribute the performance of GPT3.5 to the effectiveness of few-shot prompting using the knowledge graph. This guides GPT3.5 to create tests similar to KonTest.
GPT3.5 is slightly (4.2%) more effective than KonTest in generating tests, when provided entity relations with few-shot prompting.
Error Detection: Table 8 highlights that while 52.6% of the metamorphic errors detected by GPT3.5 were detected correctly, approximately 40.1% of the errors were false positives  [^2] Note that some errors detected by GPT3.5 cannot be automatically validated by our oracles (UNK in Table 8), as it excludes responses that do not begin with yes/no. . For ontological error detection, we provided GPT3.5 all possible queries and corresponding responses for each knowledge path. GPT3.5 was then asked to detect any inconsistency in these responses. Unlike KonTest, GPT3.5 is unable to detect multiple inconsistencies in a knowledge path. Hence, for a fair comparison in this experiment, we count ontological errors for KonTest at the granularity of a knowledge path (i.e., at most one error per knowledge path). Since validating GPT3.5 responses is not straightforward in this case, we manually validated the responses. We observed that the number of real ontological errors detected by GPT3.5 is comparable to the ontological errors detected by KonTest, but it also had a high false positive rate with nearly 63.7% of errors being misclassified.
GPT3.5 detects both metamorphic and ontological errors, but exhibits a false positive rate of up to 63.7%.
Table 7: Test generation effectiveness of GPT3.5 and KonTest on the places entities. (Meta.: Metamorphic, Onto.: Ontological)
LLM (Subject)
Valid Test
Executions (%)
Errors (%)
Meta. Onto. All Meta. Onto. All
GPT3.5 Falcon 1322 (89.9) 263 (93.9) 1585 (90.6) 232 (17.5) 25 (9.5) 257 (16.2)
Gemini 1470 (100) 280 (100) 1750 (100) 172 (11.7) 41 (14.6) 213 (12.2)
Llama2 1351 (91.9)) 264 (94.3) 1615 (92.3) 370 (27.4) 27 (10.2) 397 (24.6)
Total 4143 (93.9) 807 (96.1) 4950 (94.3) 774 (18.7) 93 (11.5) 867 (17.5)
KonTest
w/o GPT3.5
4143 (93.9) 827 (94.4) 4970 (94.0) 742 (17.9) 92 (11.1) 834 (16.8)
6 Related Work
Table 8: Error Detection Effectiveness (TP/FP: True/False positive, UNK: Unknown)
Subject Metamorphic Errors (%) Ontological Errors (%)
Kon- Test GPT3.5 Kon- Test GPT3.5
TP FP UNK Paths TP FP
Falcon 226 208 (27.7) 453 (60.4) 89 (11.9) 100 23 23 (33.3) 46 (66.7)
Gemini 161 161 (97.6) 4 (2.4) 0 (0) 100 30 30 (54.5) 25 (45.5)
Llama2 355 352 (77.0) 93 (20.4) 12 (2.6) 100 22 20 (26.0) 57 (74.0)
Total 742 721 (52.6) 550 (40.1) 101 (7.4) 300 75 73 (36.3) 128 (63.7)
LLMs and Knowledge: Pan et al. Pan et al. (2023) presents techniques to combine LLMs and knowledge graphs to address their individual limitations. As an example, Huang et al.  Huang et al. (2023c) have demonstrated the feasibility of knowledge transfer to improve LLM’s generalization ability in software engineering tasks. Analogously, WEAVER Yang et al. (2023a) uses LLMs to generate knowledge bases, using which, requirements are extracted for testing models in real-world settings. GPT4GEO Roberts et al. (2023) experimentally evaluates the capabilities and limitations of GPT-4 in geospatial domain (e.g., places entity), highlighting potential usage of GPT-4 in navigation. Unlike these proposed techniques, KonTest employs knowledge graphs to expose inconsistencies and measure knowledge gaps in LLMs.
Testing and Analysis of LLMs: Several researchers have surveyed the challenges and opportunities for testing and analysing LLMs Zhao et al. (2023); Hou et al. (2023); Zheng et al. (2023); Aleti (2023). Researchers have also studied or proposed methods for testing and analyzing several quality properties of LLMs, including their reasoning ability Wu et al. (2023); Qiu et al. (2023), non-determinism Ouyang et al. (2023), interpretability Palacio et al. (2023); Rodriguez-Cardenas et al. (2023), robustness Zhu et al. (2023), fairness Huang et al. (2023a), security concerns (e.g., privacy, memorization and backdoor attack) Yang et al. (2023b); Staab et al. (2023); Huang et al. (2023b). In contrast to the aforementioned works, KonTest studies the consistency and knowledge coverage of LLMs.
7 Conclusion
In this paper, we propose KonTest where the key intuition is to distill (a subset of) facts from a knowledge graph, which, are subsequently used to generate queries and formulate test cases for detecting a variety of consistency errors. Our evaluation reveals realistic consistency errors across state-of-the-art LLMs. Moreover, KonTest opens opportunities to investigate LLMs through the lens of their knowledge gaps, which, in turn is indicated as part of our framework. This helps designers and end users to understand and mitigate the effect of such knowledge gaps e.g., via prompt engineering or fine tuning. In future, we aim to investigate automated mitigation of consistency errors by leveraging the KonTest framework. We provide our code and experimental data in the following:
http://anonymous.4open.science/r/Kontest
8 Limitations and Threats to Validity
Construct Validity: This relates to the metrics and measures employed in our experimental analysis. To mitigate this threat, we have employed standard testing metrics such as the number/rate of generated inputs, error-inducing inputs, (knowledge) coverage and testing time. For automatic analysis of hundreds of responses, our analysis does not handle expressive, non-binary LLM responses. However, we mitigate this by employing system prompts to ensure the model provides binary responses and we discard non-binary responses (as invalid).
Internal Validity: This refers to the threat that our implementation of KonTest performs its intended knowledge graph-based test generation. We conduct several manual and automated tests, as well as inspection of sampled outcomes of KonTest to ensure the correctness of our approach. Our ablation study (RQ4) further allows to probe the correctness of the sub-step of KonTest versus using GPT3.5. We experimentally verify that over 90% of the entities we evaluate against existed in Wikidata prior to 2020. We also find that under 15% of the errors found by KonTest are linked to these entities. In addition, we also note that the SUT ought to answer the question in a consistent manner regardless of whether the entity existed prior to the corresponding knowledge cutoff date.
External Validity: The main threat to external validity of this work is the generalizability of KonTest and findings to LLMs, knowledge graphs, templates and entities beyond the ones used in this work. For instance, we mitigate the LLM generalizability threat by employing state-of-the-art off-the-shelf, mature, open model weights LLMs (LLaMA2 and FALCON), as well as commercial LLMs (such as GPT3.5 and Gemini). Notably, our entity selection and template construction may be limited to our experimental setting. However, we demonstrate the applicability of KonTest by using two different domains with multiple relations and entity types. We note that we employ few-shot prompting to query the LLMs, thus, our findings may not generalise to other prompting techniques. Finally, KonTest employs Wikidata, a well-maintained knowledge graph that is popularly used in both academia and industry Peters et al. (2019).
LLM Stability and Correctness: Researchers have identified several stability concerns about LLMs Ouyang et al. (2023); Fan et al. (2023), such as non-determinism, randomness, sensitivity to prompting methods, and API/model updates. To mitigate these threats we performed several actions: First, we set the temperature of all models to zero (0), when possible Cloud (2023); Ouyang et al. (2023). Secondly, we reduce the risk of model updates by limiting our testing time (to about a day each per model) and checking for news of model updates before and after testing. Thirdly, we also use models with frozen weights (Falcon and Llama2) to reduce the non-determinism. To automatically validate model outcomes, we employ few-shot prompting, which has been shown to be effective for querying LLMs Deng et al. (2023). We examined whether LLMs are comparable to KonTest (RQ4) using GPT3.5 since it produces the most valid responses (99.9%) (see Table 3).
Knowledge Graph Completeness and Soundness: We note that the knowledge graph is an incomplete snapshot of the real-world. In our evaluation, KonTest only tests for facts (positive tests) derived from the knowledge graph. While it is also possible to use KonTest to test for incorrect relation (negative test), the validation of such test results is challenging due to the incompleteness of knowledge graph. Moreover, we only tests 50 paths from this graph. However, these concerns do not affect our findings, as we are certain about the errors found within the tested subset of the knowledge graph. Finally, knowledge of the world naturally evolves over time and the knowledge graphs do not evolve at the same pace Pan et al. (2023). To mitigate this, we use a widely used knowledge graph Vrandečić and Krötzsch (2014).
9 Ethics Statement
We elucidate our ethics statement in this section: (1) Dataset: We utilize data from Wikidata, a publicly available open knowledge base related to Wikipedia. Wikidata is licensed under the Creative Commons CC0 License. (2) Human Evaluations: Our experiments do not involve human participants. (3) Approach: We test KonTest with the aid of LLMs (both proprietary and free). We acknowledge that these models may give biased results due to their training data and methods. However, we restrict our queries to existing relations in the knowledge graph making it unlikely to raise ethical concerns. We limit ourselves to running inference on pre-trained models due to the numerous environmental concerns (energy and water expenditure) associated with training these LLMs.
Appendix A Appendix (Additional Figures and Experimental Results)
Algorithm 3 SUT’s Knowledge Graph Generation
1:procedure Graph_Build(${LLM_{Resp}}$)
2:     ${Graph_{Gen}}\leftarrow\varnothing$
3:     for $Resp$ in ${LLM_{Resp}}$ do
4:         $\triangleright$ Finds nodes involved in response.
5:         ${Node_{s}},{Node_{d}}\leftarrow$ Find_Nodes($Resp$)
6:         if $Node_{s}\notin{Graph_{Gen}}$ then
7:              ${Graph_{Gen}}\leftarrow Graph_{Gen}\cup Node_{s}$
8:         if $Node_{d}\notin{Graph_{Gen}}$ then
9:              ${Graph_{Gen}}\leftarrow Graph_{Gen}\cup Node_{d}$
10:         if $Resp$ == Yes then
11:              ${Graph_{Gen}}\leftarrow{Graph_{Gen}}\cup$ Edge(${Node_{d}},{Node_{s}}$)
12:     return $Graph_{Gen}$
Table 9: Time performance for KonTest and GPT3.5
Approach Knowledge Construction Test Generation Error Detection
Metamorphic Ontology
KonTest Total Time 23039s 5.5ms 3s 2s
Average Time
(Per Subject)
5759.8s 1.4ms 0.8s 0.5s
GPT3.5 Total Time 345s 2331s 13232s 2711s
Average Time
(Per Subject)
115s 777s 4411s 904s
Figure 4: Example queries and templates in KonTest
Figure 5: Example queries generated by the unseen template 3 in KonTest
Figure 6: Venn Diagram of the consistency errors for each SUT.
Table 10: Effectiveness of KonTest using a knowledge graph.
LLM (Subject) Valid Test Executions (%) Errors (%)
Metamorphic Ontological All Metamorphic Ontological All
atomic
sequential
-intra
sequential
-inter
atomic
sequential
-intra
sequential
-inter
Places Falcon 254 (86.4) 530 (90.1) 538 (91.5) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 25 (9.4) 251 (15.8)
Gemini 294 (100) 588 (100) 588 (100) 292 (100) 1762 (100) 51 (17.3) 6 (1.0) 104 (17.7) 40 (13.7) 201 (11.4)
GPT3.5 293 (99.7) 588 (100) 587 (99.8) 292 (100) 1760 (99.9) 61 (20.8) 201 (34.2) 138 (23.5) 17 (5.8) 417 (23.7)
Llama2 266 (90.5) 543 (92.3) 542 (92.2) 268 (91.8) 1619 (91.9) 94 (35.3) 52 (9.6) 209 (38.6) 27 (10.1) 382 (23.6)
Total 1107 (94.1) 2249 (95.6) 2255 (95.9) 1119 (95.8) 6730 (95.5) 256 (23.1) 283 (12.6) 603 (26.7) 109 (9.7) 1251 (18.6)
Music Falcon 141 (96.6) 287 (98.3) 287 (98.3) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 3 (3.2) 134 (16.6)
Gemini 146 (100) 292 (100) 292 (100) 96 (100) 826 (100) 39 (26.7) 8 (2.7) 77 (26.4) 7 (7.3) 131 (15.9)
GPT3.5 146 (100) 292 (100) 292 (100) 96 (100) 826 (100) 40 (27.4) 85 (29.1) 56 (19.2) 5 (5.2) 186 (22.5)
Llama2 137 (93.4) 283 (96.9) 276 (94.5) 96 (100) 792 (95.9) 46 (33.6) 63 (22.3) 97 (35.1) 9 (9.4) 215 (27.1)
Total 570 (97.6) 1154 (98.8) 1147 (98.2) 382 (99.5) 3253 (98.5) 149 (26.1) 194 (16.8) 299 (26.1) 24 (6.3) 666 (20.5)
Total 1677 (95.3) 3403 (96.7) 3402 (96.6) 1501 (96.7) 9983 (96.4) 405 (24.2) 477 (14.0) 902 (26.5) 133 (8.9) 1917 (19.2)
Table 11: Metamorphic Error Detection Effectiveness of GPT3.5 (TP/FP: True/False positive, UNK: Unknown).
LLM (Subject) Error Type All
atomic sequential-intra sequential-inter
TP FP UNK TP FP UNK TP FP UNK TP FP UNK
Falcon 32 66 19 24 187 32 152 200 38 208 453 89
Gemini 51 2 0 6 2 0 104 0 0 161 4 0
llama2 94 19 5 51 42 0 207 32 7 352 93 12
Total 177 87 24 81 231 32 463 232 45 721 550 101
Table 12: Test generation effectiveness of GPT3.5 and KonTest
LLM (Subject) Valid Test Executions (%) Errors (%)
Metamorphic Ontological All Metamorphic Ontological All
atomic
sequential-intra
sequential-inter
atomic
sequential-intra
sequential-inter
Falcon 254 (86.4) 530 (90.1) 538 (91.5) 263 (93.9) 1585 (90.6) 54 (21.3) 24 (4.5) 154 (28.6) 25 (9.5) 257 (16.2)
Gemini 294 (100) 588 (100) 588 (100) 280 (100) 1750 (100) 57 (19.4) 7 (1.2) 108 (18.4) 41 (14.6) 213 (12.2)
Llama2 266 (90.5) 543 (92.3) 542 (92.2) 264 (94.3) 1615 (92.3) 97 (36.5) 61 (11.2) 212 (39.1) 27 (10.2) 397 (24.6)
Total 814 (92.3) 1661 (94.2) 1668 (94.6) 807 (96.1) 4950 (94.3) 208 (25.6) 92 (5.5) 474 (28.4) 93 (11.5) 867 (17.5)
KonTest
w/o GPT3.5
814 (92.3) 1661 (94.2) 1668 (94.6) 827 (94.4) 4970 (94.0) 195 (24.0) 82 (4.9) 465 (27.9) 92 (11.1) 834 (16.8)
Table 13: Number of tests generated by KonTest and GPT3.5 for each error type
Metamorphic Ontological Total
atomic
sequential-intra
sequential-inter
KonTest Places 294 588 588 292 1762
Music 146 292 292 96 826
GPT3.5 Places 294 588 588 280 1750
Table 14: Time taken to query each SUT in seconds.
Time Taken Falcon Gemini GPT3.5 Llama2
Atomic Queries 6305 685 221 17883
Sequential Queries 16198 1269 4925 38833
Total Time 22503 1954 5146 56716

### v2

@@
 KonTest further mitigates knowledge gaps via a weighted LLM model ensemble.
-Using four state-of-the-art LLMs (Falcon, Gemini, GPT3.5, and Llama2), we show that KonTest generates 19.2% error inducing inputs (1917 errors from 9983 test inputs).
+Using four state-of-the-art LLMs (Falcon, Gemini, GPT3.5, and Llama2), we show that KonTest generates 19.2% error inducing inputs (1917 errors from 9979 test inputs).
 It also reveals a 16.5% knowledge gap across all tested LLMs.
-KonTest’s mitigation method reduces LLM knowledge gap by 32.48%.
+A mitigation method informed by KonTest’s test suite reduces LLM knowledge gap by 32.48%.
 Our ablation study further shows that GPT3.5 is not suitable for knowledge-based consistency testing because it is only 60%-68% effective in knowledge construction.
 Knowledge-based Consistency Testing of Large Language Models
-Sai Sathiesh Rajan SUTD, Singapore sai_rajan@mymail.sutd.edu.sg                        Ezekiel Soremekun Royal Holloway, University of London ezekiel.soremekun@rhul.ac.uk
-Sudipta Chattopadhyay SUTD, Singapore sudipta_chattopadhyay@sutd.edu.sg
 1 Introduction
@@
 It first automatically extracts a set of entities and entity relationships from the knowledge graph.
-This is then used to systematically generate (semantically-equivalent) queries and create test cases for the subject LLM under test (SUT) across various settings (e.g., atomic query vs.
+This is then used to systematically generate (semantically-equivalent) yes/no queries and create test cases for the subject LLM under test (SUT) across various settings (e.g., atomic query vs.
 sequential query).
 The test case generation involves minimal, one-time effort for each type of entity relation considered (e.g., only two templates for 6730 test cases) and such templates can be reused for testing arbitrary LLMs.
-Additionally, KonTest automatically mitigates knowledge gaps by leveraging the likelihood of LLM inconsistencies to construct a weighted model ensemble.
+KonTest’s test suite can additionally be used to mitigate knowledge gaps by leveraging the likelihood of LLM inconsistencies to construct a weighted model ensemble.
 To the best of our knowledge, KonTest is the first systematic approach for automatically generating consistency tests for assessing LLMs.
@@
   4. 4.
-    KonTest automatically mitigates knowledge gaps via a weighted model ensemble via the likelihood of LLM inconsistencies.
-The mitigation technique of KonTest reduces knowledge gaps by 32.48% (section 5).
+    We propose a technique that mitigates knowledge gaps via a weighted model ensemble via the likelihood of LLM inconsistencies obtained from KonTest’s test suite.
+The mitigation technique reduces knowledge gaps by 32.48% (section 5).
   5. 5.
@@
 This enables practitioners to selectively focus on the knowledge gaps highlighted by KonTest for improving the LLM (e.g., by fine-tuning).
-Figure 3: Example queries and templates in KonTest
+Figure 3: Example queries and templates for the "Places" domain in KonTest
 […]
@@
 We note that developers can implement additional templates easily if they are interested in querying different relations.
-Moreover, creation of these templates is a one-time process and the cost incurred is minimal since the number of different types of relationships is small in comparison to the number of entities.
+Moreover, creation of these templates is a one-time process and the cost incurred is minimal (<10 LoC) since the number of different types of relationships is small in comparison to the number of entities.
+We further note that our templates are directly applicable to other domains with similar relations.
+For instance, our templates for the "Places" domain can be reused for the "Food" domain since both domains have a "contains" type relation.
 Algorithm 1 KG-Based Query Generation
@@
 […]
-KonTest then uses the relative weights to calculate the final ensemble score by first multiplying the relative weights with the results for each response and summing them.
-In particular, it considers a positive and negative responses to be one and zero respectively.
+We then compute the final ensemble score by first multiplying the relative weights with the results for each response and summing them.
+In particular, we consider positive and negative responses to be one and zero respectively.
 If the ensemble score is above the midpoint (0.5), the ensemble response for the relation is considered to be positive.
@@
 To evaluate our approach (KonTest), we pose the following research questions (RQs):
-  • RQ1 Effectiveness: How effective is KonTest in exposing consistency errors in LLMs?
+  • RQ1 Effectiveness: How effective is KonTest in exposing consistency errors in LLMs? Are KonTest results stable?
   • RQ2 Knowledge Coverage: What is the SUT’s knowledge graph coverage? How much is the SUT knowledge gap revealed by KonTest?
-  • RQ3 Knowledge Gap Mitigation: How effective is KonTest’s weighted ensemble technique in improving knowledge coverage?
+  • RQ3 Knowledge Gap Mitigation: How effective is the proposed weighted ensemble technique in improving knowledge coverage?
 How does it compare to a simple majority voting?
@@
 We choose these domains as inconsistent knowledge about places may have adversarial consequences in navigation use cases of LLMs Mapbox .
-Concurrently, inconsistencies in music domain may inaccurately capture the ownership of copyrighted materials (e.g., music albums).
+Concurrently, inconsistencies in the music domain may inaccurately capture the ownership of copyrighted materials (e.g., music albums).
 For the places domain, we take the list of countries ranked by (nominal) GDP per capita and select the top ten countries as our initial entity list IMF .
@@
 Gemini Gemini (Google) gemini-pro Unknown Dec 2023
-Knowledge Gap Mitigation: We evaluate the effectiveness of KonTest’s ensemble method by determining relative weights assigned to each SUT.
+Knowledge Gap Mitigation: We evaluate the effectiveness of an ensemble method informed by KonTest’s test suite by determining the relative weights assigned to each SUT.
 We restrict the set of relations considered for this process such that a relation that yields an invalid answer in the query generation phase for any of the SUTs is not used in subsequent computations.
@@
 Concretely, we hold out 20% of the relations for evaluation and use the consistency errors associated with the rest of the relations in the computation of the weights.
-We introduce a third template to evaluate whether KonTest’s mitigation generalizes to unseen templates.
+We introduce a third template to evaluate whether our mitigation scheme generalizes to unseen templates.
 Ablation Study: We assess whether an LLM could conceivably be used to replicate each of the sub-tasks in our approach by testing its applicability on the places domain.
@@
 All experiments were conducted on a GCP VM with an N1 series machine, eight vCPUs, 30 GB of memory and one Nvidia T4 GPU.
-Table 3: Effectiveness of KonTest using a knowledge graph. (Meta.: Metamorphic, Onto.: Ontological)
-LLM (Subject)
-Valid Test
-Executions (%)
-Errors (%)
-Meta. Onto. All Meta. Onto. All
-Places Falcon 1322 (89.9) 267 (91.4) 1589 (90.2) 226 (17.1) 25 (9.4) 251 (15.8)
-Gemini 1470 (100) 292 (100) 1762 (100) 161 (11.0) 40 (13.7) 201 (11.4)
-GPT3.5 1468 (99.9) 292 (100) 1760 (99.9) 400 (27.2) 17 (5.8) 417 (23.7)
-Llama2 1351 (91.9) 268 (91.8) 1619 (91.9) 355 (26.3) 27 (10.1) 382 (23.6)
-Total 5611 (95.4) 1119 (95.8) 6730 (95.5) 1142 (20.4) 109 (9.7) 1251 (18.6)
-Music Falcon 715 (97.9) 94 (97.9) 809 (97.9) 131 (18.3) 3 (3.2) 134 (16.6)
-Gemini 730 (100) 96 (100) 826 (100) 124 (17.0) 7 (7.3) 131 (15.9)
-GPT3.5 730 (100) 96 (100) 826 (100) 181 (24.8) 5 (5.2) 186 (22.5)
-Llama2 696 (95.3) 96 (100) 792 (95.9) 206 (29.6) 9 (9.4) 215 (27.1)
-Total 2871 (98.3) 382 (99.5) 3253 (98.5) 642 (22.4) 24 (6.3) 666 (20.5)
-Total 8482 (96.4) 1501 (96.7) 9983 (96.4) 1784 (21.0) 133 (8.9) 1917 (19.2)
 5 Evaluation Results
-Table 4: Knowledge Gap Mitigation results for KonTest’s weighted ensemble (“KonTest”), versus simple majority voting ensemble (“Majority”) and the initial average knowledge gap found in the SUTs (“SUTs”).
-Best reduction in knowledge gaps are in bold text.
-“%Impr.
-” means percentage improvement over the SUT.
-Domain Relations Knowledge Gap (%)
-Template 1 Template 2 Template 3 All (%Impr.)
-KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs
-Places 224 28 (12.5) 38 (17.0) 38.25 (17.1) 49 (21.9) 77 (34.4) 65.75 (29.4) 30 (13.4) 49 (21.9) 41.5 (18.5) 107 (26.46) 164 (-12.71) 145.5
-Music 132 39 (29.5) 60 (45.5) 48.5 (36.7) 15 (11.4) 44 (33.3) 35.75 (27.1) 24 (18.2) 55 (41.7) 44.25 (33.5) 78 (39.30) 159 (-23.74) 128.5
-Total 356 67 (18.8) 98 (27.5) 86.75 (24.4) 64 (18.0) 121 (34.0) 101.5 (28.5) 54 (15.2) 104 (29.2) 85.75 (24.1) 185 (32.48) 323 (-17.88) 274
+Table 3: Effectiveness of KonTest using a knowledge graph.
+LLM (Subject) Valid Test Executions (%) Errors (%)
+Metamorphic Ontological All Metamorphic Ontological All
+atomic
+sequential
+-intra
+sequential
+-inter
+All Types atomic
+sequential
+-intra
+sequential
+-inter
+All Types
+Places Falcon 254 (86.4) 530 (90.1) 538 (91.5) 1322 (89.9) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 226 (17.1) 25 (9.4) 251 (15.8)
+Gemini 294 (100) 588 (100) 588 (100) 1470 (100) 292 (100) 1762 (100) 51 (17.3) 6 (1.0) 104 (17.7) 161 (11.0) 40 (13.7) 201 (11.4)
+GPT3.5 293 (99.7) 588 (100) 587 (99.8) 1468 (99.9) 292 (100) 1760 (99.9) 61 (20.8) 201 (34.2) 138 (23.5) 400 (27.2) 17 (5.8) 417 (23.7)
+Llama2 266 (90.5) 543 (92.3) 542 (92.2) 1351 (91.9) 268 (91.8) 1619 (91.9) 94 (35.3) 52 (9.6) 209 (38.6) 355 (26.3) 27 (10.1) 382 (23.6)
+Total 1107 (94.1) 2249 (95.6) 2255 (95.9) 5611 (95.4) 1119 (95.8) 6730 (95.5) 256 (23.1) 283 (12.6) 603 (26.7) 1142 (20.4) 109 (9.7) 1251 (18.6)
+Music Falcon 141 (96.6) 287 (98.3) 287 (98.3) 715 (97.9) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 131 (18.3) 3 (3.2) 134 (16.6)
+Gemini 146 (100) 292 (100) 292 (100) 730 (100) 96 (100) 826 (100) 39 (26.7) 8 (2.7) 77 (26.4) 124 (17.0) 7 (7.3) 131 (15.9)
+GPT3.5 146 (100) 292 (100) 292 (100) 730 (100) 96 (100) 826 (100) 40 (27.4) 85 (29.1) 56 (19.2) 181 (24.8) 5 (5.2) 186 (22.5)
+Llama2 137 (93.4) 283 (96.9) 276 (94.5) 696 (95.3) 92 (95.8) 788 (95.4) 46 (33.6) 63 (22.3) 97 (35.1) 206 (29.6) 9 (9.8) 215 (27.3)
+Total 570 (97.6) 1154 (98.8) 1147 (98.2) 2871 (98.3) 378 (98.4) 3249 (98.3) 149 (26.1) 194 (16.8) 299 (26.1) 642 (22.4) 24 (6.3) 666 (20.5)
+Total 1677 (95.3) 3403 (96.7) 3402 (96.6) 8482 (96.4) 1497 (96.5) 9979 (96.4) 405 (24.2) 477 (14.0) 902 (26.5) 1784 (21.0) 133 (8.9) 1917 (19.2)
 RQ1 Effectiveness: We found that KonTest is effective in exposing consistency errors in LLMs: 19.2% of valid test executions result in consistency errors.
-Table 3 also shows that metamorphic errors (21.0%) are more prevalent than ontological errors (8.9%).
+Table 3 also shows that metamorphic errors (21.0% = 1784/8482) are more prevalent than ontological errors (8.9% = 133/1497).
 We attribute the lower error rate of ontological errors/inputs to the high complexity of our ontological oracle (see Table 1).
@@
 In addition, we observe that a large model size does not necessarily lead to a smaller error rate than a smaller model, e.g., GPT3.5 exhibits a metamorphic error rate of 27.2% for the places entities as compared to a metamorphic error rate of 17.1% for Falcon.
+In addition, we validate the stability of our results by repeating our experiments four additional times.
+We find that the open-source models with frozen weights (Falcon and Llama2) yielded identical results when compared to the initial experiments.
+However, the closed-source models exhibited (up to 6%) lower error rates on the subsequent iterations with Gemini exhibiting a less than one percent change in error rate.
+We attribute this to changes in the underlying models OpenAI as the additional experiments were performed approximately eight (8) months after the initial experiments.
+Furthermore, we find that the error rates for all four subsequent iterations are fairly similar indicating that the results are consistent when repeated within a short duration of time (within a few days).
+In particular, we find that Gemini results do not vary at all.
+Lastly, we also examine the stability of our results by executing both atomic and sequential queries separately.
+For instance, consider Q1 and Q6a (Table 1) and the following additional oracle: LLM(Q1)=LLM(Q6a).
+We found that only GPT3.5 exhibited inconsistent results for this oracle and further noted that the absolute number of errors found was negligible (<1%).
+These experiments demonstrate the stability of our results.
 KonTest effectively exposes consistency errors in LLMS: 19.2% of valid test executions exposed a metamorphic or ontological error in LLMs.
@@
 Total 1760 458 (26.0) 581 (33.0) 290 (16.5) 1470 (83.5)
-RQ3 Knowledge Gap Mitigation: Results show that KonTest’s mitigation technique reduces the SUT’s knowledge gap by up to 39.30%.
-Table 4 shows that KonTest reduces the knowledge gap for all templates by 32.48%.
+Table 5: Knowledge Gap Mitigation results for KonTest’s weighted ensemble (“KonTest”), versus simple majority voting ensemble (“Majority”) and the initial average knowledge gap found in the SUTs (“SUTs”).
+Best reduction in knowledge gaps are in bold text.
+“%Impr.
+” means percentage improvement over the SUT.
+Domain Relations Knowledge Gap (%)
+Template 1 Template 2 Template 3 All (%Impr.)
+KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs KonTest Majority SUTs
+Places 224 28 (12.5) 38 (17.0) 38.25 (17.1) 49 (21.9) 77 (34.4) 65.75 (29.4) 30 (13.4) 49 (21.9) 41.5 (18.5) 107 (26.46) 164 (-12.71) 145.5
+Music 132 39 (29.5) 60 (45.5) 48.5 (36.7) 15 (11.4) 44 (33.3) 35.75 (27.1) 24 (18.2) 55 (41.7) 44.25 (33.5) 78 (39.30) 159 (-23.74) 128.5
+Total 356 67 (18.8) 98 (27.5) 86.75 (24.4) 64 (18.0) 121 (34.0) 101.5 (28.5) 54 (15.2) 104 (29.2) 85.75 (24.1) 185 (32.48) 323 (-17.88) 274
+RQ3 Knowledge Gap Mitigation: Results show that our proposed mitigation technique reduces the SUT’s knowledge gap by up to 39.30%.
+Table 5 shows that our technique reduces the knowledge gap for all templates by 32.48%.
 We found that the simple majority voting ensemble worsens the SUTs’ knowledge gap by up to 23.74%.
-We also observed that the mitigation performance of KonTest generalizes to an unseen template (template three (3)).
-While the mitigation performance of KonTest is slightly better on the new template than the original templates the simple majority performs worse on an unseen template.
-These results demonstrate the generalizability of KonTest’s mitigation and the efficacy of our weighted ensemble.
-We attribute the performance of KonTest’s mitigation to the likelihood and distribution of inconsistencies: Firstly, KonTest’s weighted ensemble is 42.72% more effective than the simple majority voting.
-Secondly, the distribution of inconsistencies shows that only 1.1% (15 out of 1330) of errors are found in all four SUTs while 65.7% (874 out of 1330) of errors are unique to a single SUT.
-The ensemble method of KonTest effectively reduces the SUT’s knowledge gap by 32.48%.
+We also observed that the mitigation performance of our technique generalizes to an unseen template (template three (3)).
+While the performance of our mitigation technique is slightly better on the new template than the original templates the simple majority performs worse on an unseen template.
+On the whole, we find that our weighted ensemble is 42.72% ((323-185)/323) more effective than simple majority voting at reducing the knowledge gap.
+These results demonstrate the generalizability of our proposed mitigation technique and the efficacy of our weighted ensemble.
+We attribute the performance of our technique to the likelihood and distribution of inconsistencies discovered by KonTest (Figure 4).
+Figure 4 shows that only 1.1% (15 out of 1330) of errors are found in all four SUTs while 65.7% (874 out of 1330) of errors are unique to a single SUT.
+This suggests that an ensembling scheme, informed by the relative performance of the SUTs, reduces knowledge gap.
+The proposed ensemble method effectively reduces the SUT’s knowledge gap by 32.48%.
+Figure 4: Venn Diagram of the consistency errors for each SUT.
 RQ4 Ablation study with GPT3.5: We conduct an ablation study to investigate whether a state-of-the-art LLM (GPT3.5) is as effective as KonTest in performing its three main sub-tasks – knowledge construction, test generation and error detection.
@@
 (2023).
-These results show the importance of knowledge graphs in KonTest and suggests that LLMs are not a reliable replacement for knowledge graphs.
+These results show the importance of knowledge graphs in KonTest and suggest that LLMs are not a reliable replacement for knowledge graphs.
 GPT3.5 is an ineffective surrogate for a knowledge graph: It constructs only 68% of the ground-truth entity relations (edges).
-Test Generation: Given a few (three) query examples and the relations in a knowledge graph, GPT3.5 is slightly more effective (17.5% vs 16.8%) than KonTest in test generation (see Table 7).
-We also observe that this effectiveness persists for both the metamorphic (18.7% vs 17.9%) and the ontological oracle (11.5% vs 11.1%).
-We attribute the performance of GPT3.5 to the effectiveness of few-shot prompting using the knowledge graph.
-This guides GPT3.5 to create tests similar to KonTest.
-GPT3.5 is slightly (4.2%) more effective than KonTest in generating tests, when provided entity relations with few-shot prompting.
-Error Detection: Table 8 highlights that while 52.6% of the metamorphic errors detected by GPT3.5 were detected correctly, approximately 40.1% of the errors were false positives  [^2] Note that some errors detected by GPT3.5 cannot be automatically validated by our oracles (UNK in Table 8), as it excludes responses that do not begin with yes/no.
-.
-For ontological error detection, we provided GPT3.5 all possible queries and corresponding responses for each knowledge path.
-GPT3.5 was then asked to detect any inconsistency in these responses.
-Unlike KonTest, GPT3.5 is unable to detect multiple inconsistencies in a knowledge path.
-Hence, for a fair comparison in this experiment, we count ontological errors for KonTest at the granularity of a knowledge path (i.e., at most one error per knowledge path).
-Since validating GPT3.5 responses is not straightforward in this case, we manually validated the responses.
-We observed that the number of real ontological errors detected by GPT3.5 is comparable to the ontological errors detected by KonTest, but it also had a high false positive rate with nearly 63.7% of errors being misclassified.
-GPT3.5 detects both metamorphic and ontological errors, but exhibits a false positive rate of up to 63.7%.
 Table 7: Test generation effectiveness of GPT3.5 and KonTest on the places entities. (Meta.: Metamorphic, Onto.: Ontological)
@@
 4143 (93.9) 827 (94.4) 4970 (94.0) 742 (17.9) 92 (11.1) 834 (16.8)
+Test Generation: Given a few (three) query examples and the relations in a knowledge graph, GPT3.5 is slightly more effective (17.5% vs 16.8%) than KonTest in test generation (see Table 7).
+We also observe that this effectiveness persists for both the metamorphic (18.7% vs 17.9%) and the ontological oracle (11.5% vs 11.1%).
+We attribute the performance of GPT3.5 to the effectiveness of few-shot prompting using the knowledge graph.
+This guides GPT3.5 to create tests similar to KonTest.
+GPT3.5 is slightly (4.2%) more effective than KonTest in generating tests, when provided entity relations with few-shot prompting.
+Error Detection: Table 8 highlights that while 52.6% of the metamorphic errors detected by GPT3.5 were detected correctly, approximately 40.1% of the errors were false positives  [^2] Note that some errors detected by GPT3.5 cannot be automatically validated by our oracles (UNK in Table 8), as it excludes responses that do not begin with yes/no.
+.
+For ontological error detection, we provided GPT3.5 all possible queries and corresponding responses for each knowledge path.
+GPT3.5 was then asked to detect any inconsistency in these responses.
+Unlike KonTest, GPT3.5 is unable to detect multiple inconsistencies in a knowledge path.
+Hence, for a fair comparison, we count ontological errors for KonTest at the granularity of a knowledge path (i.e., at most one error per knowledge path).
+Since validating GPT3.5 responses is not straightforward in this case, we manually validated the responses.
+We observed that the number of real ontological errors detected by GPT3.5 is comparable to the ontological errors detected by KonTest, but it also had a high false positive rate with nearly 63.7% of errors being misclassified.
+GPT3.5 detects both metamorphic and ontological errors, but exhibits a false positive rate of up to 63.7%.
 6 Related Work
@@
 7 Conclusion
-In this paper, we propose KonTest where the key intuition is to distill (a subset of) facts from a knowledge graph, which, are subsequently used to generate queries and formulate test cases for detecting a variety of consistency errors.
+In this paper, we propose KonTest where the key intuition is to distill (a subset of) facts from a knowledge graph, which are subsequently used to generate queries and formulate test cases for detecting a variety of consistency errors.
 Our evaluation reveals realistic consistency errors across state-of-the-art LLMs.
-Moreover, KonTest opens opportunities to investigate LLMs through the lens of their knowledge gaps, which, in turn is indicated as part of our framework.
+Moreover, KonTest opens opportunities to investigate LLMs through the lens of their knowledge gaps, which, in turn, is indicated as part of our framework.
 This helps designers and end users to understand and mitigate the effect of such knowledge gaps e.g., via prompt engineering or fine tuning.
@@
 We provide our code and experimental data in the following:
-http://anonymous.4open.science/r/Kontest
+https://github.com/sparkssss/KonTest
 8 Limitations and Threats to Validity
@@
 We conduct several manual and automated tests, as well as inspection of sampled outcomes of KonTest to ensure the correctness of our approach.
-Our ablation study (RQ4) further allows to probe the correctness of the sub-step of KonTest versus using GPT3.5.
+Our ablation study (RQ4) further allows us to probe the correctness of the sub-step of KonTest versus using GPT3.5.
 We experimentally verify that over 90% of the entities we evaluate against existed in Wikidata prior to 2020.
@@
 External Validity: The main threat to external validity of this work is the generalizability of KonTest and findings to LLMs, knowledge graphs, templates and entities beyond the ones used in this work.
-For instance, we mitigate the LLM generalizability threat by employing state-of-the-art off-the-shelf, mature, open model weights LLMs (LLaMA2 and FALCON), as well as commercial LLMs (such as GPT3.5 and Gemini).
-Notably, our entity selection and template construction may be limited to our experimental setting.
+We mitigate the LLM generalizability threat by employing state-of-the-art off-the-shelf, mature, open model weights LLMs (LLaMA2 and FALCON), as well as commercial LLMs (such as GPT3.5 and Gemini).
+Similarly, our entity selection and template construction may be limited to our experimental setting.
 However, we demonstrate the applicability of KonTest by using two different domains with multiple relations and entity types.
-We note that we employ few-shot prompting to query the LLMs, thus, our findings may not generalise to other prompting techniques.
+However, we note that KonTest might not be easily adapted to information that cannot be encapsulated within a knowledge graph.
+In addition, we employ few-shot prompting to conduct our ablation study, and our findings might not generalize to other prompting techniques.
 Finally, KonTest employs Wikidata, a well-maintained knowledge graph that is popularly used in both academia and industry Peters et al.
@@
 In our evaluation, KonTest only tests for facts (positive tests) derived from the knowledge graph.
-While it is also possible to use KonTest to test for incorrect relation (negative test), the validation of such test results is challenging due to the incompleteness of knowledge graph.
-Moreover, we only tests 50 paths from this graph.
+While it is also possible to use KonTest to test for incorrect relation (negative test), the validation of such test results is challenging due to the incompleteness of the knowledge graph.
+Moreover, we only test 50 paths from this graph.
 However, these concerns do not affect our findings, as we are certain about the errors found within the tested subset of the knowledge graph.
@@
 We limit ourselves to running inference on pre-trained models due to the numerous environmental concerns (energy and water expenditure) associated with training these LLMs.
-Appendix A Appendix (Additional Figures and Experimental Results)
-Algorithm 3 SUT’s Knowledge Graph Generation
-1:procedure Graph_Build(${LLM_{Resp}}$)
-2:     ${Graph_{Gen}}\leftarrow\varnothing$
-3:     for $Resp$ in ${LLM_{Resp}}$ do
-4:         $\triangleright$ Finds nodes involved in response.
-5:         ${Node_{s}},{Node_{d}}\leftarrow$ Find_Nodes($Resp$)
-6:         if $Node_{s}\notin{Graph_{Gen}}$ then
-7:              ${Graph_{Gen}}\leftarrow Graph_{Gen}\cup Node_{s}$
-8:         if $Node_{d}\notin{Graph_{Gen}}$ then
-9:              ${Graph_{Gen}}\leftarrow Graph_{Gen}\cup Node_{d}$
-10:         if $Resp$ == Yes then
-11:              ${Graph_{Gen}}\leftarrow{Graph_{Gen}}\cup$ Edge(${Node_{d}},{Node_{s}}$)
-12:     return $Graph_{Gen}$
-Table 9: Time performance for KonTest and GPT3.5
-Approach Knowledge Construction Test Generation Error Detection
-Metamorphic Ontology
-KonTest Total Time 23039s 5.5ms 3s 2s
-Average Time
-(Per Subject)
-5759.8s 1.4ms 0.8s 0.5s
-GPT3.5 Total Time 345s 2331s 13232s 2711s
-Average Time
-(Per Subject)
-115s 777s 4411s 904s
-Figure 4: Example queries and templates in KonTest
-Figure 5: Example queries generated by the unseen template 3 in KonTest
-Figure 6: Venn Diagram of the consistency errors for each SUT.
-Table 10: Effectiveness of KonTest using a knowledge graph.
-LLM (Subject) Valid Test Executions (%) Errors (%)
-Metamorphic Ontological All Metamorphic Ontological All
-atomic
-sequential
--intra
-sequential
--inter
-atomic
-sequential
--intra
-sequential
--inter
-Places Falcon 254 (86.4) 530 (90.1) 538 (91.5) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 25 (9.4) 251 (15.8)
-Gemini 294 (100) 588 (100) 588 (100) 292 (100) 1762 (100) 51 (17.3) 6 (1.0) 104 (17.7) 40 (13.7) 201 (11.4)
-GPT3.5 293 (99.7) 588 (100) 587 (99.8) 292 (100) 1760 (99.9) 61 (20.8) 201 (34.2) 138 (23.5) 17 (5.8) 417 (23.7)
-Llama2 266 (90.5) 543 (92.3) 542 (92.2) 268 (91.8) 1619 (91.9) 94 (35.3) 52 (9.6) 209 (38.6) 27 (10.1) 382 (23.6)
-Total 1107 (94.1) 2249 (95.6) 2255 (95.9) 1119 (95.8) 6730 (95.5) 256 (23.1) 283 (12.6) 603 (26.7) 109 (9.7) 1251 (18.6)
-Music Falcon 141 (96.6) 287 (98.3) 287 (98.3) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 3 (3.2) 134 (16.6)
-Gemini 146 (100) 292 (100) 292 (100) 96 (100) 826 (100) 39 (26.7) 8 (2.7) 77 (26.4) 7 (7.3) 131 (15.9)
-GPT3.5 146 (100) 292 (100) 292 (100) 96 (100) 826 (100) 40 (27.4) 85 (29.1) 56 (19.2) 5 (5.2) 186 (22.5)
-Llama2 137 (93.4) 283 (96.9) 276 (94.5) 96 (100) 792 (95.9) 46 (33.6) 63 (22.3) 97 (35.1) 9 (9.4) 215 (27.1)
-Total 570 (97.6) 1154 (98.8) 1147 (98.2) 382 (99.5) 3253 (98.5) 149 (26.1) 194 (16.8) 299 (26.1) 24 (6.3) 666 (20.5)
-Total 1677 (95.3) 3403 (96.7) 3402 (96.6) 1501 (96.7) 9983 (96.4) 405 (24.2) 477 (14.0) 902 (26.5) 133 (8.9) 1917 (19.2)
-Table 11: Metamorphic Error Detection Effectiveness of GPT3.5 (TP/FP: True/False positive, UNK: Unknown).
-LLM (Subject) Error Type All
-atomic sequential-intra sequential-inter
-TP FP UNK TP FP UNK TP FP UNK TP FP UNK
-Falcon 32 66 19 24 187 32 152 200 38 208 453 89
-Gemini 51 2 0 6 2 0 104 0 0 161 4 0
-llama2 94 19 5 51 42 0 207 32 7 352 93 12
-Total 177 87 24 81 231 32 463 232 45 721 550 101
-Table 12: Test generation effectiveness of GPT3.5 and KonTest
-LLM (Subject) Valid Test Executions (%) Errors (%)
-Metamorphic Ontological All Metamorphic Ontological All
-atomic
-sequential-intra
-sequential-inter
-atomic
-sequential-intra
-sequential-inter
-Falcon 254 (86.4) 530 (90.1) 538 (91.5) 263 (93.9) 1585 (90.6) 54 (21.3) 24 (4.5) 154 (28.6) 25 (9.5) 257 (16.2)
-Gemini 294 (100) 588 (100) 588 (100) 280 (100) 1750 (100) 57 (19.4) 7 (1.2) 108 (18.4) 41 (14.6) 213 (12.2)
-Llama2 266 (90.5) 543 (92.3) 542 (92.2) 264 (94.3) 1615 (92.3) 97 (36.5) 61 (11.2) 212 (39.1) 27 (10.2) 397 (24.6)
-Total 814 (92.3) 1661 (94.2) 1668 (94.6) 807 (96.1) 4950 (94.3) 208 (25.6) 92 (5.5) 474 (28.4) 93 (11.5) 867 (17.5)
-KonTest
-w/o GPT3.5
-814 (92.3) 1661 (94.2) 1668 (94.6) 827 (94.4) 4970 (94.0) 195 (24.0) 82 (4.9) 465 (27.9) 92 (11.1) 834 (16.8)
-Table 13: Number of tests generated by KonTest and GPT3.5 for each error type
-Metamorphic Ontological Total
-atomic
-sequential-intra
-sequential-inter
-KonTest Places 294 588 588 292 1762
-Music 146 292 292 96 826
-GPT3.5 Places 294 588 588 280 1750
-Table 14: Time taken to query each SUT in seconds.
-Time Taken Falcon Gemini GPT3.5 Llama2
-Atomic Queries 6305 685 221 17883
-Sequential Queries 16198 1269 4925 38833
-Total Time 22503 1954 5146 56716

### v3

@@
 3.1 Knowledge Graph based Test Generation
-[…]
+Knowledge Graph Construction: KonTest allows the developer to specify the list of entities that are of interest.
+With these initial set of entities, KonTest then queries an external source of knowledge to compute, up to a certain depth, all other related entities for the considered relation.
+Developers can also control this depth, allowing them to determine the degree to which the vicinity of the initial set of entities is explored.
+Once the initial knowledge base is constructed, KonTest extracts the full path, $\mathbb{KG}$, associated with a leaf node of the knowledge base.
+This path is then used to guide the test generation process.
+For example, given the leaf entity Kinawley, as shown in Figure 1, KonTest extracts the $\mathbb{KG}$ for Kinawley as $\mathit{Kinawley}\rightarrow\mathit{Ulster}\rightarrow\mathit{Ireland}$.
 Test Generation: Given a knowledge path $\mathbb{KG}$, KonTest leverages Algorithm 1 to exhaustively generate queries relating to all possible pairs of entities present in the path.
 To this end, KonTest first finds the relation, $\mathbb{KG}_{rel}$, between a pair of nodes.
-KonTest then generates a query, ${Query}$, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (9 in Algorithm 1).
-KonTest also generates the mutated query, ${Query_{M}}$, with another template (12 in Algorithm 1).
+KonTest then generates a query, Query{Query}, using $\mathbb{KG}_{rel}$ with the aid of the template shown in Figure 3 (˜9 in Algorithm 1).
+KonTest also generates the mutated query, $Query_{M}$}, with another template (˜12 in Algorithm 1).
 These templates are dependent on the relation between the two nodes in question.
@@
 11:      $\triangleright$ Generates semantically equivalent query
-12:                  ${Query_{M}}\leftarrow$ Mut_Gen(${Query}$)
+12:      $Query_{M}$$\leftarrow$ Mut_Gen(Query{Query})
 13:      $MutQuery_{list}\leftarrow MutQuery_{list}\cup\{{Query_{M}}\}$
@@
 15:  return $Err_{count},{Coverage_{count}}$
-To check the consistency in SUT’s knowledge graph, Algorithm 2 first inspects whether a direct path exists between any pair of nodes ($Node_{1},Node_{2}$) in ${Graph_{Gen}}$ (7 in Algorithm 2).
-In the absence of such a direct path, when an indirect path exists between the same pair of nodes (11 in Algorithm 2), KonTest detects an ontological consistency error and updates the error count, ${Err_{count}}$.
+To check the consistency in SUT’s knowledge graph, Algorithm 2 first inspects whether a direct path exists between any pair of nodes ($Node_{1},Node_{2}$) in $Graph_{Gen}$} (˜7 in Algorithm 2).
+In the absence of such a direct path, when an indirect path exists between the same pair of nodes (˜11 in Algorithm 2), KonTest detects an ontological consistency error and updates the error count, $Err_{count}$}.
 Intuitively, a direct path between two nodes indicates that the SUT responded positively to a question relating the two nodes.
@@
 3.3 Knowledge Gap Mitigation
-[…]
-[…]
+We propose a weighted ensemble method that leverages the test suite generated by KonTest to mitigate the discovered knowledge gaps.
+The ensemble method uses the number of metamorphic errors found for each SUT on a per relation basis to inform its scoring system.
+A relation that induces $x$ errors is given a score of 5x5-x since each relation can maximally induce five errors.
+The cumulative SUT specific score, $\mathit{Score}_{i}$, is then found by summing the scores for each relation.
+Next, the relative weight, $\mathit{W}_{i}$, for each SUT in the set of SUTs ($\mathbb{SUT}$), is computed via the following equation:
+$\mathit{W}_{i}=\frac{\mathit{Score}_{i}}{\sum\mathit{Score}_{i}}\forall i\in\mathbb{SUT}$ (1)
 We then compute the final ensemble score by first multiplying the relative weights with the results for each response and summing them.
@@
 Table 3: Effectiveness of KonTest using a knowledge graph.
-LLM (Subject) Valid Test Executions (%) Errors (%)
+LLM
+Valid Test Executions (%) Errors (%)
 Metamorphic Ontological All Metamorphic Ontological All
@@
 All Types
-Places Falcon 254 (86.4) 530 (90.1) 538 (91.5) 1322 (89.9) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 226 (17.1) 25 (9.4) 251 (15.8)
+Places
+Falcon 254 (86.4) 530 (90.1) 538 (91.5) 1322 (89.9) 267 (91.4) 1589 (90.2) 50 (19.7) 24 (4.5) 152 (28.3) 226 (17.1) 25 (9.4) 251 (15.8)
 Gemini 294 (100) 588 (100) 588 (100) 1470 (100) 292 (100) 1762 (100) 51 (17.3) 6 (1.0) 104 (17.7) 161 (11.0) 40 (13.7) 201 (11.4)
@@
 Total 1107 (94.1) 2249 (95.6) 2255 (95.9) 5611 (95.4) 1119 (95.8) 6730 (95.5) 256 (23.1) 283 (12.6) 603 (26.7) 1142 (20.4) 109 (9.7) 1251 (18.6)
-Music Falcon 141 (96.6) 287 (98.3) 287 (98.3) 715 (97.9) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 131 (18.3) 3 (3.2) 134 (16.6)
+Music
+Falcon 141 (96.6) 287 (98.3) 287 (98.3) 715 (97.9) 94 (97.9) 809 (97.9) 24 (17.0) 38 (13.2) 69 (24.0) 131 (18.3) 3 (3.2) 134 (16.6)
 Gemini 146 (100) 292 (100) 292 (100) 730 (100) 96 (100) 826 (100) 39 (26.7) 8 (2.7) 77 (26.4) 124 (17.0) 7 (7.3) 131 (15.9)
@@
 Table 4: Knowledge Coverage and Gap in tested LLMs (Highest coverage or gap are marked in bold text)
-LLM (Subject) Knowledge Gap (%) Overall Coverage (%)
+LLM
+Knowledge Gap (%)
+Overall
+Coverage (%)
 Relations Template 1 Template 2 Intersection
-Places Falcon 294 12 (4.1) 93 (31.6) 9 (3.1) 285 (96.9)
+Places
+Falcon 294 12 (4.1) 93 (31.6) 9 (3.1) 285 (96.9)
 Gemini 294 60 (20.4) 71 (24.1) 40 (13.6) 254 (86.4)
@@
 Llama2 294 38 (12.9) 139 (47.3) 37 (12.6) 257 (87.4)
-Music Falcon 146 20 (13.7) 15 (10.3) 3 (2.1) 143 (97.9)
+Music
+Falcon 146 20 (13.7) 15 (10.3) 3 (2.1) 143 (97.9)
 Gemini 146 60 (41.1) 45 (30.8) 33 (22.6) 113 (77.4)
@@
 Table 7: Test generation effectiveness of GPT3.5 and KonTest on the places entities. (Meta.: Metamorphic, Onto.: Ontological)
-LLM (Subject)
+LLM
 Valid Test
@@
 Meta. Onto. All Meta. Onto. All
-GPT3.5 Falcon 1322 (89.9) 263 (93.9) 1585 (90.6) 232 (17.5) 25 (9.5) 257 (16.2)
+GPT3.5
+Falcon 1322 (89.9) 263 (93.9) 1585 (90.6) 232 (17.5) 25 (9.5) 257 (16.2)
 Gemini 1470 (100) 280 (100) 1750 (100) 172 (11.7) 41 (14.6) 213 (12.2)
@@
 Subject Metamorphic Errors (%) Ontological Errors (%)
-Kon- Test GPT3.5 Kon- Test GPT3.5
+Kon-
+Test
+GPT3.5
+Kon-
+Test
+GPT3.5
 TP FP UNK Paths TP FP
@@
 We limit ourselves to running inference on pre-trained models due to the numerous environmental concerns (energy and water expenditure) associated with training these LLMs.
+10 Acknowledgment
+This work was partially funded by grant number SMU-SUTD 2023_02_04 and the Singapore Ministry of Education (MOE) Present’s Graduate Fellowship.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the view of the respective funding agencies.